# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 2.17 Physics-informed neural networks
Physics-informed neural networks offer a reliable technique which combines physics concepts with the flexibility and learning capabilities of neural networks. These models tend to be useful for tackling complex physical problems when obtaining analytical solutions may be difficult or impossible to achieve. This technique can solve physics-based partial differential equations with the help of neural network architecture. It consists of two main components: a neural network and a loss function. The neural network parameterizes the solution to the problem, while the loss function ensures that the solution is close to the governing physics equations. Here, we describe a general methodology to deal with such problems. Let us consider a nonlinear partial differential equation in $\Omega\subset\mathbb{R}^{n}$ raissi2019 :
@@
where $\{x_{u}^{j},t_{u}^{j},{u}^{j}\}_{j=1}^{N_{u}}$ are the initial and boundary training data on $u(x,t)$ and $\{x_{f}^{j},t_{f}^{j}\}_{j=1}^{N_{f}}$ ate the collocations points for $f(x,t)$. If we apply this methodology to Navier-Stokes equations in two dimensions, we obtain the mean squared error (MSE) loss function as raissi2019 :
@@
where $f:=u_{t}+\lambda_{1}(uu_{x}+vu_{y})+p_{x}-\lambda_{2}(u_{xx}+u_{yy})$ and $g:=v_{t}+\lambda_{1}(uv_{x}+vv_{y})+p_{y}-\lambda_{2}(v_{xx}+v_{yy})$ with $p(x,y,t)$ as the pressure and $\lambda_{1}$ and $\lambda_{2}$ are the unknown parameters.
@@ 2.18 Modelling nonequilibrium phenomena and processes with nonlocal models
Using nonlocal models to study nonequilibrium phenomena adds an intriguing dimension to understanding dynamic processes due to the interaction over long distances. Traditional equilibrium models assume that systems are in a stable state, but many real-world situations involve constant change and flux. In the context of nonequilibrium phenomena, nonlocal models can be applied to various fields, such as physics, fluid dynamics, and materials science. Nonlocal models can describe the behaviour of particles and waves in systems that deviate from equilibrium. For instance, they can be used to study the transport of heat or charge in materials with non-uniform structures. Here, we describe a nonlocal model of an electron-hole semiconductor plasma, which accounts for the nonequilibrium phenomena. An electron system’s electro-hydrodynamic model is commonly expressed as blotekjaer1970 ; melnik2000 :
@@
where $\mathbf{z}=(n,\mathbf{v},W)^{T}$, $\zeta=(\mathcal{F}_{1},\mathbf{\mathcal{F}}_{2},\mathcal{F}_{3})^{T}$, $\mathcal{F}_{1}=-\nabla\cdot(n\mathbf{v})$, $\mathbf{\mathcal{F}}_{2}=-\mathbf{v}\cdot\nabla\mathbf{v}-q\mathbf{E}_{\mbox{eff}}/m_{n}-\nabla(nT_{n})/(m_{n}n)$, and $\mathcal{F}_{3}=-\nabla\cdot(\mathbf{v}W)-qn\mathbf{v}\cdot\mathbf{E}_{\mbox{eff}}-\nabla\cdot(\mathbf{v}nT_{n})-\nabla\cdot\mathbf{q}$. Here, $n$ is the electron concentration, $v$ is their average velocity, $W$ is the energy density, $\mathbf{E}_{\mbox{eff}}$ is the effective electric strain, $m_{n}$ is the effective electron mass, $T_{n}$ is the electron temperature given in energetic units, and $\mathbf{q}(\sim-k\nabla T_{n})$ is the heat flow. The simplified model of this can be investigated mathematically by developing drift-diffusion models; however, they are incompatible with technological improvements. In melnik2000 , the authors reduced the computational cost with the help of normalization and described a number of non-stationary physical phenomena in semiconductor devices, including carrier heating and velocity overshoot. This reduced model also accounts for the nonequilibrium and nonlocal character of electron-hole semiconductor plasma.
@@ 2.19 Nonlocal many-particle systems with AI and machine learning
Nonlocal interactions in multi-body systems extend beyond the immediate neighbours of individual bodies because multiple species evolve in an ecosystem and interact with themselves and others. For instance, species sense their surroundings to make decisions, e.g., the predator finds the prey, the prey avoids the predator, and the aggregation in swarms, called nonlocal sensing eftimie2007 ; levine2000 ; mogilner1999 . This nonlocal sensing happens on different scales, affecting populations’ spatial distributions leading to species aggregation, segregation, or complex patterns bastille2018 ; benhamou2014 ; martinez2020 ; potts2014 . Researchers have studied these kinds of biological problems mathematically by aggregation-diffusion equation giunta2021 ; carrillo2019 ; topaz2006 . For a single species model, it leads to the following form:
@@
where $m$ is a positive integer. This model can generate a non-uniform stationary pattern consisting of single or multiple patches with various shapes and sizes james2015 ; craig2016 . A generalization of the aggregation-diffusion equation (34) for multi-species populations is giunta2021 :
@@
where $u_{j}(x,t)$ are the densities of $N(\geq 1)$ populations at the spatial location $x$ and at time $t$. Here, $D_{j}(>0)$ is the diffusion coefficient of the population $j$ and $h_{jk}$ are constants that denote the attractive or repulsive tendencies (depending on its sign) of the population $j$ to the population $k$. Here, births and deaths have been neglected in the model because many animals make their spatial arrangement on shorter time scales. Authors in potts2019 studied the model (35) for $m=1$ with the top-hat kernel function $\phi$ and incorporated the detection behaviour of the populations distributed over habitat in animal ecosystems. They have explained the detection in various forms, e.g., direct observations of the individuals at a distance, indirect communication between individuals via marking the environment, and the memory of past interactions with other populations giunta2021 ; potts2019 .
@@
Nonlocal models for crowd dynamics come from the study of many-body issues, gaining influence from the Boltzmann model and growing within a hierarchy of mathematical models thieu2023 . The Boltzmann model is a core model for understanding many-body systems and offers a platform for extending notions to crowd dynamics. In addition, for the many-body problem, each individual’s behaviour impacts and is impacted by the activities of others. Traditional models frequently focus on local interactions, but nonlocal models incorporate effects throughout the entire population. Due to the complexities of real-world events, validating and calibrating nonlocal models for crowd dynamics is not an easy task.
@@
Data-driven nonlocal models in biology are an exciting intersection of computational methods and biological research srivastava2022 . These models leverage data to understand and simulate complex biological systems, considering interactions that occur over long distances. They can be applied to study phenomena like the spread of diseases, signalling pathways in cells, or ecological interactions in a community. By incorporating data into these models, researchers can gain a more comprehensive understanding of the underlying processes and make predictions based on real-world observations.
@@
Data-driven nonlocal models are effective tools for solving a variety of inverse issues in a range of scientific fields arridge2019 . When compared to conventional approaches, they enable more robust and accurate solutions to inverse issues by utilizing the knowledge found in nonlocal relationships within the data. These models are highly effective at reconstructing high-quality images from noisy or partial data in computer vision and medical imaging li2023 . They improve reconstruction accuracy by accounting for nonlocal similarities between image patches. In addition, the nonlocal technique in signal processing makes it possible to recover signals from measurements that are undersampled or distorted more effectively. The goal of various ongoing research projects is to create effective algorithms to deal with these problems. Among them is the Bayesian approach, which offers a probabilistic framework for quantifying uncertainty that combines known information with empirical data in an easy-to-use manner xie2014 . Instead of treating model parameters as fixed values, it treats them as probability distributions. The posterior distribution is then obtained by applying Bayes’ theorem to update the prior distribution of beliefs and combine it with the likelihood of observed data. Because algorithms take uncertainty into account, Bayesian models are naturally resilient to small amounts of data. This is especially useful in situations when acquiring huge datasets is difficult.
@@
AI has revolutionized all fields, including biology, and helps analyze and interpret the vast amount of biological data used to develop nonlocal models. This includes tasks such as learning, reasoning, problem-solving, perception, and language understanding. On the other hand, the generative models learn the underlying patterns and structures of the data and can then create new samples that share similar characteristics with the help of machine learning or deep learning algorithms. In addition, these models provide an emerging collection of tools with enhanced characteristics for recreating segregated and whole-brain dynamics ramezanian2022 . Furthermore, generative models in AI and learning processes are based on non-equilibrium dynamics carbone2023 .
@@
There are several types of machine learning algorithms available in the literature, which may be generally classified into three categories: supervised learning, unsupervised learning, and reinforcement learning srivastava2022 . In supervised learning, the algorithm is trained using a labelled dataset to identify the mapping function from inputs to outputs, which allows the algorithm to generate predictions or classifications on previously unknown data. In the case of unsupervised learning, the algorithm searches for patterns, correlations, or structures in the data. Agent learning algorithms are used in reinforcement learning to make decisions by interacting with the environment. The agent receives feedback in the form of benefits based on the behaviours it takes. Here, we describe a data-driven technique using an optimization-based methodology to learn nonlocal constitutive laws for stress wave propagation models. For this, we consider a high-fidelity model in a domain $\Omega\subset\mathbb{R}^{d}$ you2021 :
@@
$\mathcal{L}_{K}[u]=f(x),\leavevmode\nobreak\ x\in\Omega_{I},$
@@
where $||\cdot||_{\mathcal{X}}$ an appropriate norm over $\Omega$. Authors in you2021 described a detailed algorithm to optimize this problem. In addition, this type of methodology can also be applied to hyperbolic problems.
@@ 2.20 Multiscale and multifidelity models
In computational science and engineering, various multilevel methods have been applied to mathematical models, such as multigrid for solving systems of equations, multilevel Monte Carlo and multilevel stochastic collocation for estimating mean solutions of partial differential equations with stochastic parameters, and so on. The common approach is the use of surrogate models, which are simplified models trained on a combination of high and low-fidelity data. These surrogate models can then be used to quickly approximate solutions to the nonlocal models. In these techniques, a parameter is involved and controls the trade-off between error and computational costs. Here, we first describe the Monte Carlo method for a general problem to determine approximate statistical information of the high-fidelity model with low-fidelity surrogate models peherstorfer2016 .
Let $f:\mathcal{D}\rightarrow\mathcal{Y}$ be a function for the source of information, where $\mathcal{D}\subset\mathbb{R}^{d}$ for some $d\in\mathbb{N}$ and $\mathcal{Y}\subset\mathbb{R}$. Here, we consider the high-fidelity model $f^{(1)}$ as the “truth” model and the low-fidelity models $f^{(2)},\ldots,f^{(k)}$ as surrogate models. For $j=1,\ldots,k$, we assume the costs of evaluating each of the models $f^{(j)}$ are $w_{j}\geq 0$. For the Monte Carlo method, the authors in peherstorfer2016 draw independent and identically distributed (i.i.d) realizations $\mathbf{z}_{1},\ldots,\mathbf{z}_{m}\in\mathcal{D}$ of the random variable $Z:\Omega\rightarrow\mathcal{D}$, where $\Omega$ is a sample space, and estimated $\mathbb{E}[f^{(j)}(Z)]$ by
@@
The authors in peherstorfer2016 showed the MSE for the multifidelity Monte Carlo (MFMC) method is lower than the Monte Carlo estimator with the same computational cost. The authors considered $k$ number of i.i.d realizations $\mathbf{z}_{m_{1}},\ldots,\mathbf{z}_{m_{k}}\in\mathcal{D}$ for the $k$ models $f^{(1)},\ldots,f^{(k)}$, where $m_{j}\in\mathbb{N}$ and they satisfy $0<m_{1}\leq\ldots\leq m_{k}$. In this case, the MFMC estimates
@@
where $\alpha_{2},\ldots,\alpha_{k}\in\mathbb{R}$ are the weights for the differences of the Monte Carlo estimators $\overline{y}_{m_{j}}^{(j)}$ and $\overline{y}_{m_{j-1}}^{(j)}$, $j=2,\ldots,k$. This MFMC method can be applied to get a lower cost without compromising accuracy for the steady-state nonlocal diffusion model khodabakhshi2021 :

### v2

@@ 2.16 Physics-informed neural networks
Physics-informed neural networks offer a reliable technique which combines physics concepts with the flexibility and learning capabilities of neural networks. These models are useful for tackling complex physical problems when obtaining analytical solutions may be difficult or impossible. This technique can solve physics-based PDEs with the help of neural network architecture. It consists of two main components: a neural network and a loss function. The neural network parameterizes the solution to the problem, while the loss function ensures that the solution is close to the governing physics equations. Here, we describe a general methodology to deal with such problems. Let us consider a nonlinear partial differential equation in $\Omega\subset\mathbb{R}^{n}$ [287]:
@@
where $\{x_{u}^{j},t_{u}^{j},{u}^{j}\}_{j=1}^{N_{u}}$ are the initial and boundary training data on $u(x,t)$ and $\{x_{f}^{j},t_{f}^{j}\}_{j=1}^{N_{f}}$ ate the collocations points for $f(x,t)$. If we apply this methodology to Navier-Stokes equations in two dimensions, we obtain the mean squared error (MSE) loss function as [287]:
@@
where $f:=u_{t}+\lambda_{1}(uu_{x}+vu_{y})+p_{x}-\lambda_{2}(u_{xx}+u_{yy})$ and $g:=v_{t}+\lambda_{1}(uv_{x}+vv_{y})+p_{y}-\lambda_{2}(v_{xx}+v_{yy})$ with $p(x,y,t)$ as the pressure and $\lambda_{1}$ and $\lambda_{2}$ are the unknown parameters. A fundamental advantage of utilizing neural networks in PINNs is their capacity to execute automated differentiation, which allows the network to effectively compute derivatives of the anticipated function (in terms of time and space). In addition, the key concept behind PINN techniques is to encode the underlying physical rule (i.e., the PDE) as prior information in the neural network. However, the PINN technique fails to produce realistic approximations of the solution when the PDE contains non-convex flow functions, such as shocks and mixed waves [288]. Nevertheless, training PINNs requires careful initialization, especially when working with stiff equations or highly nonlinear systems.
@@ 2.17 Modelling nonequilibrium phenomena and processes with nonlocal models
Nonequilibrium phenomena occur when a system is not in equilibrium, i.e., it has net flows of energy, matter, or information [289]. These systems often exhibit complex, time-dependent behaviour, whereas the macroscopic properties of equilibrium systems are stable and time-invariant. In addition, the entropy is constantly produced for nonequilibrium processes; however, it is maximal and constant for the equilibrium case. These nonequilibrium systems have some key characteristics, such as self-organization behaviours [290] (e.g., pattern formation in chemical reactions and ecological systems), irreversibility (e.g., processes such as diffusion, heat conduction, and chemical reactions get the system closer to equilibrium over time but never entirely achieve it), etc.
Using nonlocal models to investigate nonequilibrium events adds an exciting dimension to understanding dynamic processes caused by long-distance interactions. Traditional equilibrium models assume that systems are in a stable state, but many real-world situations involve constant change and flux. In the context of nonequilibrium phenomena, nonlocal models can be applied to various fields, such as physics, fluid dynamics, and materials science. Nonlocal models can describe the behaviour of particles and waves in systems that deviate from equilibrium. For instance, they can be used to study the transport of heat or charge in materials with non-uniform structures. Here, we describe a nonlocal model of an electron-hole semiconductor plasma, which accounts for the nonequilibrium phenomena. An electron system’s electro-hydrodynamic model is commonly expressed as [291, 292]:
@@
where $\mathbf{z}=(n,\mathbf{v},W)^{T}$, $\zeta=(\mathcal{F}_{1},\mathbf{\mathcal{F}}_{2},\mathcal{F}_{3})^{T}$, $\mathcal{F}_{1}=-\nabla\cdot(n\mathbf{v})$, $\mathbf{\mathcal{F}}_{2}=-\mathbf{v}\cdot\nabla\mathbf{v}-q\mathbf{E}_{\mbox{eff}}/m_{n}-\nabla(nT_{n})/(m_{n}n)$, and $\mathcal{F}_{3}=-\nabla\cdot(\mathbf{v}W)-qn\mathbf{v}\cdot\mathbf{E}_{\mbox{eff}}-\nabla\cdot(\mathbf{v}nT_{n})-\nabla\cdot\mathbf{q}$. Here, $n$ is the electron concentration, $v$ is their average velocity, $W$ is the energy density, $\mathbf{E}_{\mbox{eff}}$ is the effective electric strain, $m_{n}$ is the effective electron mass, $T_{n}$ is the electron temperature given in energetic units, and $\mathbf{q}(\sim-k\nabla T_{n})$ is the heat flow. The simplified model of this can be investigated mathematically by developing drift-diffusion models; however, they are incompatible with technological improvements. In [292], the authors reduced the computational cost with the help of normalization and described several non-stationary physical phenomena in semiconductor devices, including carrier heating and velocity overshoot. This reduced model also accounts for the nonequilibrium and nonlocal character of electron-hole semiconductor plasma.
@@ 2.18 Nonlocal many-particle systems with AI and machine learning
Nonlocal interactions in multi-body systems extend beyond the immediate neighbours of individual bodies because multiple species evolve in an ecosystem and interact with themselves and others. For instance, species sense their surroundings to make decisions, e.g., the predator finds the prey, the prey avoids the predator, and the aggregation in swarms, called nonlocal sensing [48, 49, 293]. This nonlocal sensing happens on different scales, affecting populations’ spatial distributions leading to species aggregation, segregation, or complex patterns [294, 295, 296, 297, 298, 299, 56]. Researchers have studied these kinds of biological problems mathematically by aggregation-diffusion equation [300, 301, 302, 56]. For a single species model, it leads to the following form:
@@
where $m$ is a positive integer. This model can generate a non-uniform stationary pattern consisting of single or multiple patches with various shapes and sizes [303, 304]. A generalization of the aggregation-diffusion equation (39) for multi-species populations is [300]:
@@
where $u_{j}(x,t)$ are the densities of $N(\geq 1)$ populations at the spatial location $x$ and at time $t$. Here, $D_{j}(>0)$ is the diffusion coefficient of the population $j$ and $h_{jk}$ are constants that denote the attractive or repulsive tendencies (depending on its sign) of the population $j$ to the population $k$. Here, births and deaths have been neglected in the model because many animals make their spatial arrangement on shorter time scales. Authors in [305] studied the model (40) for $m=1$ with the top-hat kernel function $\phi$ and incorporated the detection behaviour of the populations distributed over habitat in animal ecosystems. They have explained the detection in various forms, e.g., direct observations of the individuals at a distance, indirect communication between individuals via marking the environment, and the memory of past interactions with other populations [300, 305, 306].
@@
Nonlocal models for crowd dynamics come from studying many-body issues, gaining influence from the Boltzmann model and growing within a hierarchy of mathematical models [307]. The Boltzmann model is a core model for understanding many-body systems and offers a platform for extending notions to crowd dynamics. In addition, for the many-body problem, each individual’s behaviour impacts and is impacted by the activities of others [308]. Traditional models frequently focus on local interactions, but nonlocal models incorporate effects throughout the entire population. Due to the complexities of real-world events, validating and calibrating nonlocal models for crowd dynamics is not easy.
@@
Data-driven nonlocal models in biology are an exciting intersection of computational methods and biological research [309]. These models leverage data to understand and simulate complex biological systems, considering interactions that occur over long distances. They can be applied to study phenomena like the spread of diseases, signalling pathways in cells, or ecological interactions in a community. By incorporating data into these models, researchers can gain a more comprehensive understanding of the underlying processes and make predictions based on real-world observations.
@@
Data-driven nonlocal models are effective tools for solving various inverse issues in a range of scientific fields [310]. When compared to conventional approaches, they enable more robust and accurate solutions to inverse issues by utilizing the knowledge found in nonlocal relationships within the data. These models effectively reconstruct high-quality images from noisy or partial data in computer vision and medical imaging [311]. They improve reconstruction accuracy by accounting for nonlocal similarities between image patches. In addition, the nonlocal technique in signal processing makes it possible to recover signals from undersampled or distorted measurements more effectively. The goal of various ongoing research projects is to create effective algorithms to deal with these problems. Among them is the Bayesian approach, which offers a probabilistic framework for quantifying uncertainty that combines known information with empirical data in an easy-to-use manner [312]. Instead of treating model parameters as fixed values, it treats them as probability distributions. The posterior distribution is then obtained by applying Bayes’ theorem to update the prior distribution of beliefs and combine it with the likelihood of observed data. Because algorithms take uncertainty into account, Bayesian models are naturally resilient to small amounts of data. This is especially useful in situations when acquiring huge datasets is difficult.
@@
AI has revolutionized all fields, including biology [313], and helps analyze and interpret the vast amount of biological data used to develop nonlocal models. This includes tasks such as learning, reasoning, problem-solving, perception, and language understanding. On the other hand, the generative models learn the underlying patterns and structures of the data and can then create new samples that share similar characteristics with the help of machine learning or deep learning algorithms [314, 315, 316]. In addition, these models provide an emerging collection of tools with enhanced characteristics for recreating segregated and whole-brain dynamics [317]. Furthermore, generative models in AI and learning processes are based on non-equilibrium dynamics [318].
@@
Several types of machine learning algorithms are available in the literature, which can be generally classified into three categories: supervised learning, unsupervised learning, and reinforcement learning [309]. In supervised learning, the algorithm is trained using a labelled dataset to identify the mapping function from inputs to outputs, which allows the algorithm to generate predictions or classifications on previously unknown data. In the case of unsupervised learning, the algorithm searches for patterns, correlations, or structures in the data. Agent learning algorithms are used in reinforcement learning to make decisions by interacting with the environment. The agent receives feedback in the form of benefits based on the behaviours it takes. Here, we describe a data-driven technique using an optimization-based methodology to learn nonlocal constitutive laws for stress wave propagation models. For this, we consider a high-fidelity model in a domain $\Omega\subset\mathbb{R}^{d}$ [319]:
@@
${}\mathcal{L}_{K}[u]=f(x),\leavevmode\nobreak\ x\in\Omega_{I},$ (41)
@@
where $||\cdot||_{\mathcal{X}}$ an appropriate norm over $\Omega$. Authors in [319] described a detailed algorithm to optimize this problem. In addition, this type of methodology can also be applied to hyperbolic problems.
@@ 2.19 Multiscale and multifidelity models
In computational science and engineering, various multilevel methods have been applied to mathematical models, such as multigrid for solving systems of equations, multilevel Monte Carlo and multilevel stochastic collocation for estimating mean solutions of partial differential equations with stochastic parameters, and so on. The common approach is using surrogate models, simplified models trained on a combination of high and low-fidelity data. These surrogate models can then be used to quickly approximate solutions to the nonlocal models. In these techniques, a parameter is involved and controls the trade-off between error and computational costs. Here, we first describe the Monte Carlo method for a general problem to determine approximate statistical information of the high-fidelity model with low-fidelity surrogate models [320].
Let $f:\mathcal{D}\rightarrow\mathcal{Y}$ be a function for the source of information, where $\mathcal{D}\subset\mathbb{R}^{d}$ for some $d\in\mathbb{N}$ and $\mathcal{Y}\subset\mathbb{R}$. Here, we consider the high-fidelity model $f^{(1)}$ as the “truth” model and the low-fidelity models $f^{(2)},\ldots,f^{(k)}$ as surrogate models. For $j=1,\ldots,k$, we assume the costs of evaluating each of the models $f^{(j)}$ are $w_{j}\geq 0$. For the Monte Carlo method, the authors in [320] draw independent and identically distributed (i.i.d) realizations $\mathbf{z}_{1},\ldots,\mathbf{z}_{m}\in\mathcal{D}$ of the random variable $Z:\Omega\rightarrow\mathcal{D}$, where $\Omega$ is a sample space, and estimated $\mathbb{E}[f^{(j)}(Z)]$ by
@@
The authors in [320] showed the MSE for the multifidelity Monte Carlo (MFMC) method is lower than the Monte Carlo estimator with the same computational cost. The authors considered $k$ number of i.i.d realizations $\mathbf{z}_{m_{1}},\ldots,\mathbf{z}_{m_{k}}\in\mathcal{D}$ for the $k$ models $f^{(1)},\ldots,f^{(k)}$, where $m_{j}\in\mathbb{N}$ and they satisfy $0<m_{1}\leq\ldots\leq m_{k}$. In this case, the MFMC estimates
@@
where $\alpha_{2},\ldots,\alpha_{k}\in\mathbb{R}$ are the weights for the differences of the Monte Carlo estimators $\overline{y}_{m_{j}}^{(j)}$ and $\overline{y}_{m_{j-1}}^{(j)}$, $j=2,\ldots,k$. This MFMC method can be applied to get a lower cost without compromising accuracy for the steady-state nonlocal diffusion model [125]: